# Scaling Laws for Upcycling Mixture-of-Experts Language Models

### v1

@@ 1 Introduction
Large-scale neural network architectures, such as dense transformers (Vaswani et al., 2017), have seen remarkable success across a wide range of tasks, particularly achieving human-level capabilities in natural language processing (Achiam et al., 2023). However, they often demand an enormous amount of computation, imposing challenges of computational efficiency and scalability. Sparse models like mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., 2020), have emerged as an alternative achieving better efficiency-performance trade-off via partial activation (routing) of neural parameters when processing the input. Even so, MoE still requires substantial compute power to reach its full potential (Wei et al., 2024; Dai et al., 2024; Yang et al., 2024).
One direction to further accelerate training convergence is leveraging smaller pretrained models to guide the training of larger MoE models. Komatsuzaki et al. (2022) propose upcycling, which reuses the dense checkpoint to continued pretrain the upcycled MoE. The MoE is expected to specialize and optimize routing rapidly by leveraging the pretrained dense model weights.
Despite the promise of efficient MoE training via upcycling, the effectiveness and limitations of this technique remain unclear. While some (Wei et al., 2024; He et al., 2024) have already adopted it to training large-scale MoEs, Muennighoff et al. (2024) reported negative results where upcycling can slow down training convergence. We believe these seemingly contradictory conclusions are due to insufficient comprehensive studies and assessments. There is also a lack of guidance on how and when to upcycle MoE, hampering a wider adoption of this technique.
In this paper, we seek to better understand large language model’s (LLM) upcycling to MoE via a series of controlled experiments, spanning up to a few hundred billion (B) training tokens and models up to 7B total parameters. Specifically, we uncover precise power-law scalings for the language modeling performance (cross-entropy loss) with respect to training dataset size (for both dense and upcycled MoE training), and the model configuration, including the total number of parameters (model size). Building on these results, we provide a framework for assessing when upcycling offers advantages over from-scratch training and how performance gains depend on dataset size and model configuration.
@@ 2.1 Model details
Mixture-of-Experts. The MoE model in consideration is the same as our dense model, but with all MLP blocks replaced by multiple blocks (experts) with the same configuration (Fedus et al., 2022). A router consisting of a single-layer MLP outputs the routing probability of the tokens to the experts.
@@
Upcycling. The upcycling scenario assumes that one is given a pretrained dense model and would like to train an MoE with the same configuration but replacing the MLPs with the MoE modules (Komatsuzaki et al., 2022). By replicating the dense model’s MLP weight $n_{\rm expert}$ to form the experts, the knowledge from the dense model can be reused to accelerate the training of the MoE compared to training the MoE from scratch (from random parameter initialization). Other modules’ weights are copied from the dense counterparts directly, with the router’s initial weights randomized. See Figure 2 for an illustration of upcycling. We employ this technique for our study.
@@ 2.2 Power-law Ansatz
There is an extensive literature showing that the loss of training deep learning models has a simple power-law behavior: $L=\frac{A}{X^{\alpha}}+E$, for single variable $X$, including model size and dataset size (Hestness et al., 2017, 2019; Rosenfeld et al., 2019; Henighan et al., 2020).  [^1] The actual form we assume is $L=\frac{A}{(X+1)^{\alpha}}+E$ such that the loss is finite at the limit $X\to 0$. However, since the values of $X$ we consider are often much larger than 1 ($10^{6}$ or more), we approximate it as $L=\frac{A}{X^{\alpha}}+E$ for notational convenience. We use this ansatz in this work.
@@ 3.1 Setup
Dense models in consideration. We train a suite of dense models with model sizes 15M, 44M, 0.1B, 0.2B, 0.5B and 1B. The model configuration (number of layers, $n_{\rm layer}$, hidden dimension $d_{\rm model}$, and MLP hidden dimension, $d_{\rm mlp}$) used in this paper is summarized in Table LABEL:tab:config of Appendix A.
@@
Dataset. We use training dataset derived from the CommonCrawl portion of Slimpajama-DC (Shen et al., 2023), containing 368B tokens in total. The test loss is calculated from the default validation set (0.3B tokens) defined therein. In Appendix B, we train models on two different datasets (Japanese language and source code datasets) to study the scaling behavior across datasets.
@@ 4.2 Scaling Law for Model Configuration
Using these variables, we hypothesize, similar to Equation 1, the following scaling relation:
$L(P,N_{2})=BP^{-\beta_{1}}N_{2}^{-\beta_{2}+\beta_{3}\log P}+E$ (11)
To fit the above equation, we run experiments fixing $D_{1}=D_{2}=2.3\text{B}$, varying the following: dense model of size {15M, 44M, 0.1B, 0.2B}, $n_{\rm expert}=\{4,8,16,32\}$ and $n_{\rm TopK}=\{1,2,4\}$. The scatter plot is shown in the left panel of Figure 1.
We obtain the following values for the exponents using the same fitting procedure in Section 4.2: $\beta_{1}=0.03\;,\beta_{2}=0.30\;,\beta_{3}=-0.01$. See Figure 5 for the fits. Unlike dataset sizes, due to computational constraints, we are unable to generate more data points to make a more precise estimate extending the range of model configuration. Nevertheless, we do not observe performance trade-offs when scaling sparsity and active parameters within the range of our examination, in contrast to the scaling law of dataset sizes. Our work suggests that: Performance of upcycling improves with sparsity and active parameters without noticable trade-offs.
@@ 5 Joint Scaling Law
Requirement 4. ${\rm lim}_{D_{2}\to 0}L(D_{1},D_{2},N_{1})=AD_{1}^{-\alpha}+BN_{1}^{-\beta}+E$, i.e., approaching the Chinchilla scaling law.
The straightforward functional form that satisfies these requirements is as follows.  [^3] One could extend this with Equation 11 for a truly unified scaling law including all model configurations, but we avoid doing so in this work as reliably fitting with that many independent variables would require extensive experiments beyond our computational reach.
$L(D_{1},D_{2},N_{2})=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+BN_{1}^{-\beta_{2}}+E$ (12)
@@
In Figure 6, we show that the fitted result of upcycling scaling law extrapolates well, achieving validation RMS error of 0.015. The fitted parameters using all data for all three scaling laws are summarized in Table 2. We consider other fitting possibilities and present more results in Appendix D. For the rest of this Section, we consider the application and implication of the joint scaling law.
@@ 5.1 Training MoE from Scratch versus Upcycling
Past studies (Komatsuzaki et al., 2022; Muennighoff et al., 2024) have explored the efficiency of training an MoE from scratch compared to upcycling a dense model. Komatsuzaki et al. (2022) found that upcycling retains an advantage up to 120% of the sunk cost ($D_{1}$). For instance, to match the performance of an upcycled MoE trained with an additional 0.4 trillion (T) tokens after 2T dense tokens, training an MoE from scratch would require 2.4T tokens—effectively saving 2T tokens in upcycled training. Conversely, Muennighoff et al. (2024) reported that training from scratch requires less than 100% of the sunk cost under different settings, indicating that it could be more efficient. These seemingly contradictory results suggest that upcycling efficiency depends on both sunk cost and model size.
@@ 5.2 Compute Allocation and Compute-Optimal Upcycling
We find that for a given compute budget without any pretrained models, training a compute-optimal dense or MoE model from scratch outperforms the two-stage dense-to-upcycled MoE approach (see Figure 15 in Appendix E.2). This suggests: If no pretrained model exists, direct training is preferable to upcycling for optimal performance.
In scenarios where a pretrained dense model is already available, we can determine the compute-optimal scaling of MoE upcycling. The compute cost of upcycled training is approximated as $C_{2}=6N_{2}D_{2}$ (Kaplan et al., 2020). Optimizing the compute leads to the scaling relations (see Appendix E.2 for derivation):
@@ 6 Related Work
Mixture-of experts. While interests in developing open MoE LLMs are relatively recent (Hu et al., 2024; Yang et al., 2024; Dai et al., 2024; Liu et al., 2024; Sun et al., 2024), the use of MoE in deep learning can be traced to early 2010s (Eigen et al., 2013; Bengio et al., 2013). See Cai et al. (2024) for a detailed survey of modern MoE models.
Upcycling. Leveraging pretrained models to expedite the training of larger dense models is also known as model growth (Chen et al., 2015). In the context of training MoEs reusing dense pretrained models, Komatsuzaki et al. (2022) are the first studying such a scenario with encode-decoder models. There are studies (Hu et al., 2024; Yang et al., 2024; Lo et al., 2024; Wei et al., 2024; Muennighoff et al., 2024; He et al., 2024) offering insights into upcycling decoder-only transformers, systematic investigation has not been presented. Wei et al. (2024) made only a rough recommendation: use upcycling when the budget for training is smaller than twice the budget used for dense training, while we provide a more general guideline.
Scaling laws. Power-law scaling appears in a variety of natural and human-made phenomena. Scaling studies that are perhaps closest to our work are those considering two-stage training, e.g., transfer learning, fine-tuning, and model growth (Mikami et al., 2022; Zhang et al., 2024a; Du et al., 2024). A notable similar phenomenon in transfer learning is ossification, where pretraining can hurt fine-tuning performance (Hernandez et al., 2021). However, to our knowledge, no prior work incorporates an interaction term ($\alpha_{3}$ in Equation 1) to capture such effects.
@@ 7 Discussion and Conclusion
While our empirical formulae successfully capture the observed scaling behavior, the underlying mechanism, particularly the interaction term, remains theoretically unexplained. To our knowledge, formal justification for such interactions is lacking in the literature (Bahri et al., 2024; Paquette et al., 2024). Future work could explore it through optimization dynamics or loss landscape analysis. Additionally, extending our scaling laws to alternative MoE architectures and encoder-decoder models would help assess their generality and implications.
@@ A.1 Megatron-LM configuration
Infrastructure. Our experiments are performed on multiple nodes, each consisting of 8 NVIDIA H100 80 GB GPUs, interconnected via InfiniBand HDR. The software we use for training is the Megatron-LM library [Shoeybi et al., 2019].
We use and modify the Megatron-LM (core v0.8.0) library for our experiments.  [^4] https://github.com/NVIDIA/Megatron-LM Models are trained with data type bfloat16. Except for the largest MoE we train (8x1B), which has tensor parallelism configured to be 2, all models are trained with data and sequence parallelisms only [Korthikanti et al., 2023]. Other optimization libraries used include FlashAttention [Dao et al., 2022] and TransformerEngine.  [^5] https://github.com/NVIDIA/TransformerEngine See the example scripts included in the supplementary material.
@@ A.2 Model configuration
Let us elaborate more on our architectural choices. The intermediate hidden dimension size, $d_{\rm MLP}$, is set to be $4d_{\rm model}$. We do not implement bias in the linear layers. We also do not use techniques geared for treating training instabilities (which we did not encounter in our study), such as Z-loss or QK normalization. Efficiency-motivated implementations like grouped query attention are not considered as well for simplicity. The number of attention head is chosen to increase with model size following practices in the LLM literature. Other designs of the architecture follow Llama2’s closely [Touvron et al., 2023]. See Table LABEL:tab:config for the model configurations. They are selected such that the ratio $n_{\rm layer}/d_{\rm model}$ lies in the range 32 to 64, as in Kaplan et al. [2020]. We use the smaller models for ablation studies.
@@ A.3 MoE configuration
Note that this is also known as token-choice algorithm. Furthermore, we do not use the token-dropping mechanism as in Fedus et al. [2022]. We also do not study MoE variants such as shared experts [Dai et al., 2024] and fine-grained experts [Krajewski et al., 2024], as upcycling these variants is not straightforward.
Let us move to discussing the load-balancing loss. It has the form [Fedus et al., 2022]:
@@ A.5 Training configuration
The common setup of training is shown in Table 4, and the model-dependent setup (warmup iteration, standard deviation of the normal distribution for initializing weights, maximum iteration run, battch size, tuned LR) is shown in Table 5. As described in the main text, we use the WSD schedule for training. The number of warmup steps of the WSD LR schedule is set to be roughly the same as the total model size [Porian et al., 2024]. Linear decay to 10% of the maximum LR value is used in the last stage of the schedule, with the length set to be around 10% of the training length, following Hägele et al. [2024].
Logarithmically-spaced intermediate checkpoints are saved and used to emulate different numbers of token budget. We also increase both the training length and batch size with model size following common practices without performing precise tuning.
@@
Optimizer AdamW [Loshchilov et al., 2017]

### v2

@@ Abstract
language modeling, mixture of experts, scaling law, upcycling
@@ 1 Introduction
Large-scale neural network architectures, such as dense transformers (Vaswani et al., 2017), have seen remarkable success across a wide range of tasks, particularly achieving human-level capabilities in natural language processing (Achiam et al., 2023). However, they often demand an enormous amount of computation, imposing challenges of computational efficiency and scalability. Sparse models like mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., ) have emerged as an alternative achieving better efficiency-performance trade-off via partial activation (routing) of neural parameters when processing the input. Even so, MoE models still require substantial compute power to reach its full potential (Wei et al., 2024; Dai et al., 2024; Yang et al., 2024).
One direction to further accelerate training convergence is leveraging smaller pretrained models to guide the training of larger MoE models. Komatsuzaki et al. proposed upcycling, which reuses the dense checkpoint to continued pretrain the upcycled MoE. The MoE is expected to specialize and optimize routing more rapidly by leveraging the pretrained dense model weights.
Despite the promise of efficient MoE training via upcycling, the effectiveness and limitations of this technique remain unclear. While some (Wei et al., 2024; He et al., 2024) have already adopted it to training large-scale MoE models, Muennighoff et al. (2024) reported negative results where upcycling can slow down training convergence. We believe these seemingly contradictory conclusions are due to insufficient comprehensive studies and assessments. There is also a lack of guidance on how and when to upcycle, hampering a wider adoption of this technique.
In this paper, we seek to better understand large language models’ (LLMs) upcycling to MoE models via a series of controlled experiments, spanning up to a few hundred billion (B) training tokens and models up to 7B total parameters. Specifically, we uncover precise power-law scalings for the language modeling performance (cross-entropy loss) with respect to training dataset size (for both dense and upcycled MoE training), and the model configuration, including the total number of parameters (model size). Building on these results, we provide a framework for assessing when upcycling offers advantages over from-scratch training and how performance gains depend on dataset size and model configuration.
@@ 2.1 Model details
Mixture-of-Experts. The MoE in consideration is the same as our dense model, but with all MLP blocks replaced by multiple blocks (experts) with the same configuration (Fedus et al., 2022). A router consisting of a single-layer MLP outputs the routing probability of the tokens to the experts.
@@
Upcycling. The upcycling scenario assumes that one is given a pretrained dense model and would like to train an MoE with the same configuration but replacing the MLPs with the MoE modules (Komatsuzaki et al., ). By replicating the dense model’s MLP weights $n_{\rm expert}$ times to form the experts, the knowledge from the dense model can be reused to accelerate the training of the MoE compared to training the MoE from scratch (from random parameter initialization). Other modules’ weights are copied from the dense counterparts directly, with the router’s initial weights randomized. See Figure 2 for an illustration of upcycling. We employ this technique for our study. See Appendix A.6 for other details, including alternative upcycling methods.
@@ 2.2 Power-law Ansatz
There is an extensive literature showing that the loss of training deep learning models has a simple power-law behavior: $L=\frac{A}{X^{\alpha}}+E$, for single variable $X$, including model size and dataset size (Hestness et al., 2017, 2019; Rosenfeld et al., ; Henighan et al., 2020) [^1] The actual form we assume is $L=\frac{A}{(X+1)^{\alpha}}+E$ such that the loss is finite at the limit $X\to 0$. However, since the values of $X$ we consider are often much larger than 1 ($10^{6}$ or more), we approximate it as $L=\frac{A}{X^{\alpha}}+E$ for notational convenience.. We use this ansatz in this work.
@@ 3.1 Setup
Dense models in consideration. We train a suite of dense models with model sizes 15 million (M), 44M, 0.1B, 0.2B, 0.5B and 1B. The model configuration (number of layers, $n_{\rm layer}$, hidden dimension $d_{\rm model}$, and MLP hidden dimension, $d_{\rm mlp}$) used in this paper is summarized in Table LABEL:tab:config of Appendix A.
@@
Dataset. We use training dataset derived from the CommonCrawl portion of Slimpajama-DC (Shen et al., 2023), containing 368B tokens in total. The test loss is calculated from the default validation set (0.3B tokens) defined therein. In Appendix B, we train models on two different datasets (Japanese language and source code datasets) to show that the scaling behavior generalizes across datasets.
@@ 4.2 Scaling Law for Model Configuration
Using these variables, we can derive and validate the appropriate functional form for the scaling law, similar to the analysis performed in Section 4.1. We fit the functional form by running experiments fixing $D_{1}=D_{2}=2.3\text{B}$, varying the following: dense model of size {15M, 44M, 0.1B, 0.2B}, $n_{\rm expert}=\{4,8,16,32\}$ and $n_{\rm TopK}=\{1,2,4\}$. The scatter plot is shown in the left panel of Figure 1 (for better readability, see the 2D slices of the plot with fixed sparsity in Figure 13). To this end, we find that the following functional form provides the best fit to the data:
$L(P,N_{2})=AP^{-\beta_{1}}+FN_{2}^{-\beta_{2}}+E$ (11)
See Appendix D.1 for details.
We obtain the following values for the exponents: $\beta_{1}=1.87\;,\beta_{2}=0.34$. See Figure 5 for the fits. Unlike dataset sizes, due to computational constraints, we are unable to generate more data points to make a more precise estimate extending the range of model configuration. Nevertheless, we do not observe performance trade-offs when scaling sparsity and active parameters within the range of our examination, in contrast to the scaling law of dataset sizes. Our work suggests that: Performance of upcycling improves with sparsity and active parameters without noticable trade-offs.
@@ 5 Joint Scaling Law
Requirement 4. ${\rm lim}_{D_{2}\to 0}L(D_{1},D_{2},N_{1})=AD_{1}^{-\alpha}+BN_{1}^{-\beta}+E$, i.e., approaching Chinchilla scaling law. The straightforward functional form that satisfies these requirements is as follows  [^3] One could extend this with Equation 11 to create a truly unified scaling law encompassing all model configurations. However, we refrain from doing so in this work, as reliably fitting such a scaling law with many independent variables would require significantly more extensive experimentation beyond our current computational capabilities..
$L(D_{1},D_{2},N_{1})=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+BN_{1}^{-\beta_{2}}+E$ (12)
@@
In Figure 6, we show that the fitted result of upcycling scaling law extrapolates well, achieving validation RMS error of 0.015. The fitted parameters using all data for all three scaling laws are summarized in Table 2. We consider other fitting possibilities and present more results in Appendix E. For the rest of this Section, we consider the application and implication of the joint scaling law.
@@ 5.1 Training MoE from Scratch versus Upcycling
Past studies (Komatsuzaki et al., ; Muennighoff et al., 2024) have explored the efficiency of training an MoE from scratch compared to upcycling a dense model. Komatsuzaki et al. found that upcycling retains an advantage up to 120% of the sunk cost ($D_{1}$). For instance, to match the performance of an upcycled MoE trained with an additional 0.4 trillion (T) tokens after 2T dense tokens, training an MoE from scratch would require 2.4T tokens—effectively saving 2T tokens in upcycled training. Conversely, Muennighoff et al. (2024) reported that training from scratch requires less than 100% of the sunk cost under different settings, indicating that it could be more efficient. These seemingly contradictory results suggest that upcycling efficiency depends on both sunk cost and model size.
@@ 5.2 Compute Allocation and Compute-Optimal Upcycling
We find that for a given compute budget without any pretrained models, training a compute-optimal dense or MoE from scratch outperforms the two-stage dense-to-upcycled MoE approach (see Figure 16 in Appendix F.2). This suggests: If no pretrained model exists, direct training is preferable to upcycling for optimal performance.
In scenarios where a pretrained dense model is already available, we can determine the compute-optimal scaling of MoE upcycling. The compute cost of upcycled training is approximated as $C_{2}=6N_{2}D_{2}$ (Kaplan et al., 2020). Optimizing the compute leads to the scaling relations (see Appendix F.2 for derivation):
@@ 6 Related Work
Mixture-of-Experts. While interests in developing open MoE LLMs are relatively recent (Hu et al., 2024; Yang et al., 2024; Dai et al., 2024; Liu et al., 2024; Sun et al., 2024), the use of MoE in deep learning can be traced to early 2010s (Eigen et al., 2013; Bengio et al., 2013). See Cai et al. (2024) for a detailed survey of modern MoE models.
Upcycling. Leveraging pretrained models to expedite the training of larger dense models is also known as model growth (Chen et al., 2015). In the context of training MoEs reusing dense pretrained models, Komatsuzaki et al. are the first studying such a scenario with encode-decoder models. There are studies (Hu et al., 2024; Yang et al., 2024; Lo et al., 2024; Wei et al., 2024; Muennighoff et al., 2024; He et al., 2024) offering insights into upcycling decoder-only transformers. However, a systematic investigation has been lacking. Wei et al. (2024) made only a rough recommendation: use upcycling when the budget for training is smaller than twice the budget used for dense training. In contrast, we have presented a more general guideline.
Scaling laws. Power-law scaling appears in a variety of natural and human-made phenomena. Scaling studies that are perhaps closest to our work are those considering two-stage training, e.g., transfer learning, fine-tuning, and model growth (Mikami et al., 2022; Zhang et al., ; Du et al., 2024). Prior work on transfer learning has proposed scaling laws using multiplicative or additive forms involving $D_{1}$ (pretraining data) and $D_{2}$ (fine-tuning data) (Mikami et al., 2022; Zhang et al., ). A notable similar phenomenon in transfer learning is ossification, where pretraining can hurt fine-tuning performance (Hernandez et al., 2021). However, to our knowledge, no prior work incorporates an interaction term ($\alpha_{3}$ in Equation 1) to capture such effects. Although our experiments specifically focus on upcycling into MoE models, we believe that the core insights, such as the interaction between dense and upcycled training budgets, are useful for formulating two-stage training regimes more broadly, including transfer learning as mentioned above, and potentially model growth.
@@ 7 Discussion and Conclusion
While our empirical formulae successfully capture the observed scaling behavior, the underlying mechanism, particularly the interaction term, remains theoretically unexplained. To our knowledge, formal justification for such a term is lacking in the literature (Paquette et al., ). Future work could explore it through, e.g., simplified models (Hutter, 2021; Maloney et al., 2022; Lin et al., ; Bordelon et al., 2024; Bahri et al., 2024; Paquette et al., ). Additionally, extending our scaling laws to alternative MoE architectures (shared experts and fine-grained experts (Dai et al., 2024; Krajewski et al., 2024)), modalities (vision transformer (Zhai et al., 2022)), and other multi-stage training paradigms (as discussed in Section 6) would help assess their generality and implications.
Software and Data
The source code and data (cross-entropy losses) for analyses of the paper is available at https://github.com/sbintuitions/sparse-upcycling-scaling-laws.
Impact Statement
This paper works toward the goal of advancing the field of machine learning and language modeling, with an emphasis on scaling. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Megatron-LM configuration
Infrastructure. Our experiments are performed on multiple nodes, each consisting of 8 NVIDIA H100 80 GB GPUs, interconnected via InfiniBand HDR. The software we use for training is the Megatron-LM library (Shoeybi et al., 2019).
We use and modify the Megatron-LM (core v0.8.0) library for our experiments [^4] https://github.com/NVIDIA/Megatron-LM. Models are trained with data type bfloat16. Except for the largest MoE we train (8x1B), which has tensor parallelism configured to be 2, all models are trained with data and sequence parallelisms only (Korthikanti et al., 2023). Other optimization libraries used include FlashAttention (Dao et al., 2022) and TransformerEngine [^5] https://github.com/NVIDIA/TransformerEngine. See the example scripts provided on Github666dense model example: https://github.com/NVIDIA/Megatron-LM/tree/main/examples/gpt3; MoE example: https://github.com/NVIDIA/Megatron-LM/tree/main/examples/mixtral.
@@ A.2 Model configuration
Let us elaborate more on our architectural choices. The intermediate hidden dimension size, $d_{\rm MLP}$, is set to be $4d_{\rm model}$. We do not implement bias in the linear layers. We also do not use techniques geared for treating training instabilities (which we did not encounter in our study), such as Z-loss or QK normalization. Efficiency-motivated implementations like grouped query attention are not considered as well for simplicity. The number of attention head is chosen to increase with model size following practices in the LLM literature. Other designs of the architecture follow Llama2’s closely (Touvron et al., 2023). See Table LABEL:tab:config for the model configurations. They are selected such that the ratio $n_{\rm layer}/d_{\rm model}$ lies in the range 32 to 64, as in Kaplan et al. (2020). We use the smaller models for ablation studies.
@@ A.3 MoE configuration
Note that this is also known as token-choice algorithm. Furthermore, we do not use the token-dropping mechanism as in Fedus et al. (2022). We also do not study MoE variants such as shared experts and fine-grained experts (Dai et al., 2024; Krajewski et al., 2024), as upcycling these variants is not straightforward.
Let us move to discussing the load-balancing loss. It has the form (Fedus et al., 2022):
@@ A.5 Training configuration
The common setup of training is shown in Table 4, and the model-dependent setup (warmup iteration, standard deviation of the normal distribution for initializing weights, maximum iteration run, battch size, tuned LR) is shown in Table 5. As described in the main text, we use the WSD schedule for training. The number of warmup steps of the WSD LR schedule is set to be roughly the same as the total model size (Porian et al., 2024). Linear decay to 10% of the maximum LR value is used in the last stage of the schedule, with the length set to be around 10% of the training length, following Hägele et al. (2024).
Logarithmically-spaced intermediate checkpoints are saved and used to emulate different numbers of training token budget. We also increase both the training length and batch size with model size following common practices without performing precise tuning.
@@
Optimizer AdamW (Loshchilov et al., 2017)